OpenStudy (anonymous):

Given the following definition, compute Q(5). Q(n) = 0 if n = 0 1 if n = 1 3 if n = 2 Q(n − 1) + Q(n − 2) + Q(n − 3) if n > 2

OpenStudy (anonymous):

Q(0) = 0 Q(1) = 1 Q(2) = 2 Q(3) = Q(2) + Q(1) + Q(0) = 2 + 1 + 0 = 3 Q(4) = Q(3) + Q(2) + ............... Get the idea?

OpenStudy (anonymous):

You have to calculate all the Q() s up through Q(4) in order to calculate Q(5). I've already done most of it for you.

OpenStudy (anonymous):

how did you get q(0)?

OpenStudy (anonymous):

Right from your question. You wrote: Q(n) = 0 if n=0 That means Q(0) = 0